OpenStudy (anonymous):

Can someone please explain how to find the inverse of sin^(−1)(sin(−7pi/4)) ?

OpenStudy (anonymous):

find an angle between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) with the same sine

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

I thought so, but since it is a negative input, would I have to subract -7pi/4 from 2pi ?

OpenStudy (psymon):

Hmm....since no-one has answered your last question. If you havea negative angle, you can add 2pi to it to get an equivalent positive angle. -7pi/4 + 2pi = pi/4

OpenStudy (anonymous):

I've tried pi/4 but its still wrong

OpenStudy (psymon):

I wish I could tell you why then. I agree with satellite, pi/4 should be a correct response. Of course I could be missingsomething.

OpenStudy (anonymous):

Okay well thanks for trying at least.

OpenStudy (anonymous):

:/